OpenStudy (anonymous):

simplifying: ((3t)/(a+t+2)(a+2))/t

OpenStudy (anonymous):

look up simplify expression calculator it will help you with this

OpenStudy (anonymous):

I know the steps just don't know why

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

like i know you cancel the t's out

OpenStudy (anonymous):

but wouldn't you be left with 2t at the top?

OpenStudy (anonymous):

oh nevermind

OpenStudy (anonymous):

I get it

OpenStudy (anonymous):

\[ \frac{\frac{a}{b}}{c} = \frac{a}{bc} \]

OpenStudy (jhannybean):

\[\large \frac{\frac{3t}{(a+t+2)(a+2)}}{t}\]\[=\frac{3t}{(a+t+2)(a+2)}\cdot \frac{1}{t}\]

OpenStudy (skullpatrol):